OpenStudy (anonymous):

For what value of a and b , the lines represented by equation ax^2 + xy + by^2=0 , will be inclined at 45 degrees?

OpenStudy (shubhamsrg):

by^2 + yx + ax^2 =0 =>y = -x + sqrt(x^2 - 4abx^2)/2 and -x -2 sqrt(x^2 - 4abx^2)/2 or y= x(-1 + sqrt(1 - 4ab)/2 and x(-1 - sqrt(1 - 4ab)/2 these are the eqns of the 2 lines let m1 = (-1 + sqrt(1 - 4ab)/2 and m2 = (-1 - sqrt(1 - 4ab)/2 since the 2 lines are at 45 degrees to each other,we have 1 = | m1- m2 / 1+ m1 m2 | obtain2 eqns, solve for a and b.

OpenStudy (anonymous):

how can we equate slopes to lines?

OpenStudy (shubhamsrg):

didnt get your query ?

OpenStudy (anonymous):

what are m1 and m2?

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

but we will get only 1 eq.

OpenStudy (shubhamsrg):

its in mod right,,so you'll get 2 eqns.. like this if |x| = 1 then x= 1 and -1..

OpenStudy (anonymous):

ok